OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

y-5 = 2(x+1) y-5 = 2x+2 y-7 = 2x 2x-y = -7

OpenStudy (anonymous):

thanks:))